Rational Equations. Solve the equation.
(8x-8)/(x^2-4)=4/(x-2)

Answer by Susmitha(21)

(8x-8)/(x^2-4)=4/(x-2)
(8x-8)(x-2)=4(x^2-4)
8x^2-16x-8x+16=4x^2-16
4x^2-24x+32=0
x^2-6x+8=0
x^2-4x-2x+8=0
x(x-4)-2(x-4)=0
(x-4)(x-2)=0
x=4,x=2